OpenStudy (yisrael012332):

What is 2 to the power of 3 over 2 equal to?

OpenStudy (yisrael012332):

@dan815

OpenStudy (yisrael012332):

@Michele_Laino

OpenStudy (yisrael012332):

@nincompoop

OpenStudy (yisrael012332):

@nono266

OpenStudy (anonymous):

what do u think

OpenStudy (yisrael012332):

3 sqr root 16

OpenStudy (yisrael012332):

What is 2 to the power of 3 over 2 equal to? square root of 8 cube root of 8 cube root of 16 square root of 16

OpenStudy (yisrael012332):

so what do you think

OpenStudy (yisrael012332):

anyone?

OpenStudy (anonymous):

the answer is D 2^3 equals 8 and when you divide 8 by 2 you get 4 and the square root of 16 is 4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

sqrt of 16

OpenStudy (yisrael012332):

thanks so much guys

imqwerty (imqwerty):

5487453409/5487453409 x 4 is the answer

OpenStudy (michele_laino):

hint: \[\Large {2^{3/2}} = \sqrt {{2^3}} = ...?\]